A motor bike starting from rest accelerates at a constant rate of 3.0ms-2
What is the velocity of the car after it travelled 45.5m
1
Expert's answer
2020-05-08T16:06:19-0400

We know that:

"x=x_0+v_0\times t+ \dfrac{a\times t^2}{2}"

Or:

"l=x-x_0=v_0\times t+ \dfrac{a\times t^2}{2}"

For this task "v_0=0" , so:

"l=\dfrac{a\times t^2}{2} => t=\sqrt{\dfrac{2l}{a}}=\sqrt{\dfrac{2*45.5}{3}}=\sqrt{\dfrac{91}{3}}\approx5.51 s(seconds)"

We know that:

"a=\dfrac{v-v_0}{t}=\dfrac{v}{t}=> v=a\times t= 3\times 5.51=16.53 mps" (meters per second)